OpenStudy (anonymous):

How much less would a 23 year old female pay for a $25,000 policy of 20 year life insurance (@ $2.90 per $1000) than straight life (@ $15.78 per $1000)

OpenStudy (marziajackson$):

What are the choices?

OpenStudy (marziajackson$):

the answer is $322.00

OpenStudy (anonymous):

Rate * coverage amount / Premium Scale 20 year- 2.90 * 25,000 / 1,000 = 72.50 straight life policy- 15.78 * 25,000 / 1,000 = 394.50 394.50 - 72.50 = 322.00